# Quasi-boolean groups

### v2

@@ 2. $(ii)$
Now let $c\in H^{n}_{\mathfrak{A}}(G,\mathbb{Z}/p)$ be arbitrary, and let $N\triangleleft G$ be an open normal subgroup such that there is a
$\underline{d}=\{d_{A}\in H^{\bullet}(A,\mathbb{Z}/p)\mid A\in\mathfrak{A}(G,N)\}\in\underline{H}^{n}_{\mathfrak{A}}(G,N,\mathbb{Z}/p)$
with the property $c=\mathfrak{a}_{N}(\underline{d})$. By Lemma 7.22, there is an open, normal subgroup $M\triangleleft G$ such that $M\subseteq N$ and $\mathfrak{A}(G,M,N)=\mathfrak{A}(G,N)$. For every $A\in\mathfrak{A}(G/M)$, let $\pi_{M,N,A}\colon A\to\pi_{M,N}(A)$ be the map induced by the restriction of $\pi_{M,N}$ onto $A$. The fact $\mathfrak{A}(G,M,N)=\mathfrak{A}(G,N)$ means that the collection
$e=\left\{\pi^{*}_{M,N,A}(d_{\pi_{M,N}(A)})\in H^{\bullet}(A,\mathbb{Z}/p)\mid A\in\mathfrak{A}(G/M)\right\}$
is a well-defined element of $H^{n}_{\mathfrak{A}}(G/M,\mathbb{Z}/p)=\underline{H}^{n}_{\mathfrak{A}}(G/M,\mathbb{Z}/p)$. Applying Quillen’s theorem to $e$ we get that there is a positive integer $m$ and an $f\in H^{n}(G/M,\mathbb{Z}/p)$ such that $e^{p^{m}}=q_{G/M}(f)$. Let $\phi_{M}\colon\underline{H}^{n}_{\mathfrak{A}}(G/M,\mathbb{Z}/p)\to\underline{H}^{n}_{\mathfrak{A}}(G,M,\mathbb{Z}/p)$ be the map induced by the inclusion functor $\mathfrak{A}(G,M)\to\mathfrak{A}(G/M)$ as above. Then $\mathfrak{a}_{M}(\phi_{M}(e))=\mathfrak{a}_{N}(\underline{d})$, and hence
@@
$\pi_{i}\colon H^{i}(G,\mathbb{Z}/2)\to C(\mathcal{X}^{*}(G),\mathbb{Z}/2)$
@@
Let $C^{\bullet}(\mathcal{X}^{*}(G),\mathbb{Z}/2)$ denote the graded Boolean algebra associated to the Boolean ring $C(\mathcal{X}^{*}(G),\mathbb{Z}/2)$. By Proposition 7.11, the conditions of Proposition 7.17 apply to $G$. Hence, by Theorem 7.20, there is an $F$-isomorphism:
$\pi_{*}\colon H^{\bullet}(G,\mathbb{Z}/2)\to C^{\bullet}(\mathcal{X}^{*}(G),\mathbb{Z}/2).$
This map is zero on the nilradical of $H^{\bullet}(G,\mathbb{Z}/2)$, so it induces an $F$-isomorphism:
(7.23.1) $H^{\bullet}(G,\mathbb{Z}/2)/\mathcal{N}(H^{\bullet}(G,\mathbb{Z}/2))\to C^{\bullet}(\mathcal{X}^{*}(G),\mathbb{Z}/2)$
of graded Boolean algebras by Lemma 7.7, which must be an isomorphism by Lemma 7.13. Since the nilradical of $H^{\bullet}(G,\mathbb{Z}/2)$ consists of degree one elements, the claim follows. ∎
@@ 8. Proof of the main theorem and some consequences
The implication $(i)\Rightarrow(ii)$ is Theorem 6.6, while the implication $(ii)\Rightarrow(i)$ is Theorem 6.17. We already saw that $(ii)$ trivially implies $(iii)$ in the proof of Theorem 5.4. Now assume that $G$ is the maximal pro-$2$ quotient of a real projective profinite group $H$. As explained in [8, Section 10], $H^{\bullet}(H,\mathbb{Z}/2)$ is the connected sum of a dual algebra and a graded Boolean algebra. The pull-back map $H^{\bullet}(H,\mathbb{Z}/2)\to H^{\bullet}(G,\mathbb{Z}/2)$ is an isomorphism by the Rost–Voevodsky norm residue theorem [14], so we get that $G$ is cohomologically quasi-Boolean. Therefore the implication $(iii)\Rightarrow(iv)$ holds. To finish the proof of Theorem 1.6, we show the remaining implication $(iv)\Rightarrow(i)$ in the following
@@
Since $G$ has an open subgroup without $2$-torsion by part $(i)$ of Proposition 7.11, it will be sufficient to show that every real embedding problem for $G$ has a solution. By Proposition 5.2, we need to show that any real $2$-embedding problem for $G$ has a solution. In fact we will prove something stronger.
@@
For every group $G$, let $G^{\#}$ denote the set of its conjugacy classes, for every $x\in G$, let $x^{\#}\in G^{\#}$ denote the conjugacy class of $x$, and, for every homomorphism $h\colon G\to H$ of groups, let $h^{\#}\colon G^{\#}\to H^{\#}$ denote the map on conjugacy classes induced by $h$. Now let $G$ be a profinite group. An embedding problem with lifting data $(\mathbf{E},f)$ for $G$ is an embedding problem $\mathbf{E}$:
@@
$\chi^{*}(o(\mathbf{E}))=o(\mathbf{E}(\chi)),$
where $\chi^{*}\colon H^{\bullet}(G,\mathrm{Ker}(\mathbf{E}))\to H^{\bullet}(H,\mathrm{Ker}(\mathbf{E}))$ is the pull-back map.
@@
Let $G$ be a quasi-Boolean pro-$2$ group $G$. Then every $2$-embedding problem with lifting data for $G$ and with a kernel isomorphic to $\mathbb{Z}/2$ has a solution.
@@
Let $(\mathbf{E},f)$ be an embedding problem with lifting data for $G$ as in Definition 8.2, and assume that its kernel is isomorphic to $\mathbb{Z}/2$. Since the automorphism group of the latter is trivial, we get that $\mathbf{E}$ is central. Because $\mathbf{E}$ is equipped with lifting data, it is real, that is, for every subgroup $H\subseteq G$ of order $2$ the embedding problem $\mathbf{E}(i_{H})$ has a solution, where $i_{H}\colon H\to G$ is the inclusion map. Therefore, by Remark 8.5, the image of $o(\mathbf{E})$ under the homomorphism:
$\pi_{2}\colon H^{2}(G,\mathbb{Z}/2)\to C(\mathcal{X}^{*}(G),\mathbb{Z}/2)$
@@
Let $s\colon G\to B$ be such a solution. Let $r\colon\mathcal{X}^{*}(G)\to\mathbb{Z}/2=\mathrm{Ker}(\alpha)$ be the map given by the rule:
@@
Since $s^{\#}\times f\colon\mathcal{X}^{*}(G)\to B^{\#}\times B^{\#}$ is continuous with finite image, and $r(x)$ only depends on $s^{\#}(x)$ and $f(x)$ for every $x\in\mathcal{X}^{*}(G)$, we get that $r$ is continuous. Therefore, by Corollary 7.23, there is a continuous homomorphism $\chi\colon G\to\mathbb{Z}/2=\mathrm{Ker}(\alpha)$ whose image of under the homomorphism:
$\pi_{1}\colon H^{1}(G,\mathbb{Z}/2)\to C(\mathcal{X}^{*}(G),\mathbb{Z}/2)$
@@
using that $s,\chi$ are homomorphisms and $\mathrm{Ker}(\alpha)$ is central. Therefore $\widetilde{\phi}$ is a homomorphism. Since $\alpha\circ\widetilde{\phi}=\alpha\circ s=\phi$, we get that $\widetilde{\phi}$ is a solution to $\mathbf{E}$. Now let $y\in\mathcal{Y}^{*}(G)$ be arbitrary. If $s^{\#}(y^{\#})=f(y^{\#})$, then $\widetilde{\phi}(y)=s(y)$, and hence $\widetilde{\phi}(y)^{\#}=s(y)^{\#}=s^{\#}(y^{\#})=f(y^{\#})$. If $s^{\#}(y^{\#})\neq f(y^{\#})$, then $\widetilde{\phi}(y)\neq s(y)$, so $\widetilde{\phi}(y)$ is the unique element of $\alpha^{-1}(\phi(y))$ different from $s(y)$. Since $\alpha^{-1}(\phi(y))$ contains an element of $f(y^{\#})$, as $\alpha$ is surjective, and this element is not $s(y)$, it must be $\widetilde{\phi}(y)$. So $\widetilde{\phi}(y)^{\#}=f(y^{\#})$ in this case, too. ∎
@@
such that the kernel of the quotient map $\pi_{k}\colon B/N_{k}\to B/N_{k+1}$ is isomorphic to $\mathbb{Z}/2$ for every $k=0,1,\ldots,n-1$. Let $q_{k}\colon B\to B/N_{k}$ be the quotient map. Note that it will be sufficient to show that for every continuous homomorphism $h:G\to B/N_{k+1}$ such that $h^{\#}|_{\mathcal{X}^{*}(G)}=q_{k+1}^{\#}\circ f=\pi_{k}^{\#}\circ q_{k}^{\#}\circ f$ the embedding problem $\mathbf{E}_{k}$:
@@
with lifting data $q_{k}^{\#}\circ f$ has a solution for every $k=0,1,\ldots,n$. Indeed let $r_{k}\colon B/N_{k}\to A=B/N_{n}$ be the quotient map; then we would get by descending induction on the index $k$ that the embedding problem:
@@
with lifting data $q_{k}^{\#}\circ f$ has a solution. The claim is now clear from the case $k=0$. However $\mathbf{E}_{k}$ has a kernel isomorphic to $\mathbb{Z}/2$, so $(\mathbf{E}_{k},q_{k}^{\#}\circ f)$ has a solution by Lemma 8.6. ∎
@@
First assume that $G$ is Boolean, i.e., $G$ is isomorphic to $\mathbb{B}(X)$ for some profinite space $X$. Then it is cohomologically quasi-Boolean by Theorem 1.6, so $H^{\bullet}(G,\mathbb{Z}/2)$ is the connected sum of a dual algebra $D^{\bullet}$ and a graded Boolean algebra $B^{\bullet}$. Assume that $D^{\bullet}$ is non-trivial, so there there is a non-zero $c\in D^{1}\subset\textrm{Hom}(G,\mathbb{Z}/2)$. Then $c\cup c=0$ as $D^{2}=0$, so the restriction of $c$ onto every involution in the image of $i_{X}\colon X\to\mathcal{Y}(\mathbb{B}(X))$ is zero. Therefore, by the universal property of $\mathbb{B}(X)$ of Remark 6.3, the homomorphism corresponding to $c$ is also zero, a contradiction, and hence $G$ is cohomologically Boolean.
Next assume that $G$ is cohomologically Boolean. Then it is quasi-Boolean by Theorem 1.6, so $G$ is the free product of a free pro-$2$ group $F$ and a Boolean group $\mathbb{B}(X)$ for a profinite space $X$. Assume that $F$ is non-trivial; then there is a non-zero $c\in\textrm{Hom}(F,\mathbb{Z}/2)$. Let $\pi_{1}\colon G\to F$ be the surjective homomorphism supplied by the universal property of free pro-$2$ products, and let $\overline{c}$ be the composition of $\pi_{1}$ and $c$. Then $c\cup c=0$, and hence $\overline{c}\cup\overline{c}=0$. Since $H^{\bullet}(G,\mathbb{Z}/2)$ is a graded Boolean algebra, this implies that $\overline{c}$ is zero, a contradiction. Therefore $G$ is Boolean. ∎
@@
Let $G$ be a Boolean group and let $f\colon G\to\mathbb{Z}/2$ be a quasi-canonical homomorphism. By Remark 7.9, the homomorphism $f$ is characterised by the property that, for every $n>0$ and $c\in H^{n}(G,\mathbb{Z}/2)$, we have $c^{2}=c\cup k^{n}$. But $H^{\bullet}(G,\mathbb{Z}/2)$ is a graded Boolean algebra by Corollary 1.7, and hence $f$ is unique. Therefore, it is justified to call it the canonical homomorphism of the Boolean group $G$.
@@
By assumption, $G$ is isomorphic to $F(Y^{\prime})*_{2}\mathbb{B}(X^{\prime})$ for a set $Y^{\prime}$ and a profinite space $X^{\prime}$. We are actually going to show something stronger, namely, that the profinite spaces $X$ and $X^{\prime}$ are homeomorphic, and the sets $Y^{\prime}$ and $Y$ are bijective. By Theorem 6.11, the profinite spaces $X^{\prime}$ and $\mathcal{X}^{*}(G)$ are homeomorphic, and, as we saw in the proof of Corollary 7.23, the graded Boolean algebras $B^{\bullet}$ and $C^{\bullet}(\mathcal{X}^{*}(G),\mathbb{Z}/2)$ are isomorphic (see (7.23.1)). Therefore, by Stone duality Theorem 2.10, the profinite spaces $X$ and $X^{\prime}$ are homeomorphic, too. Using Notation 6.16, we have $G_{*}=F(Y^{\prime})_{*}\oplus\mathbb{B}(X^{\prime})_{*}$. Hence we have $H^{1}(G,\mathbb{Z}/2)=H^{1}(F(Y^{\prime}),\mathbb{Z}/2)\oplus H^{1}(\mathbb{B}(X^{\prime}),\mathbb{Z}/2)$. But $H^{1}(G,\mathbb{Z}/2)=D^{1}\oplus B^{1}$, as since $H^{1}(\mathbb{B}(X^{\prime}),\mathbb{Z}/2)\cong B^{1}$ by the above, we get that $H^{1}(F(Y^{\prime}),\mathbb{Z}/2)\cong D^{1}$. Therefore, $\mathbb{Z}/2^{\oplus Y^{\prime}}\cong\mathbb{Z}/2^{\oplus Y}$, and hence there is a bijection between $Y^{\prime}$ and $Y$. ∎

### v3

@@ 2. $(ii)$
Now let $c\in H^{n}_{\mathfrak{A}}(G,\mathbb{Z}/p\mathbb{Z})$ be arbitrary, and let $N\triangleleft G$ be an open normal subgroup such that there is a
$\underline{d}=\{d_{A}\in H^{\bullet}(A,\mathbb{Z}/p\mathbb{Z})\mid A\in\mathfrak{A}(G,N)\}\in\underline{H}^{n}_{\mathfrak{A}}(G,N,\mathbb{Z}/p\mathbb{Z})$
with the property $c=\mathfrak{a}_{N}(\underline{d})$. By Lemma 7.23, there is an open normal subgroup $M\triangleleft G$ such that $M\subseteq N$ and $\mathfrak{A}(G,M,N)=\mathfrak{A}(G,N)$. For every $A\in\mathfrak{A}(G/M)$, let $\pi_{M,N,A}\colon A\to\pi_{M,N}(A)$ be the map induced by the restriction of $\pi_{M,N}$ onto $A$. The fact $\mathfrak{A}(G,M,N)=\mathfrak{A}(G,N)$ means that the collection
$e=\left\{\pi^{\bullet}_{M,N,A}(d_{\pi_{M,N}(A)})\in H^{\bullet}(A,\mathbb{Z}/p\mathbb{Z})\mid A\in\mathfrak{A}(G/M)\right\}$
is a well-defined element of $H^{n}_{\mathfrak{A}}(G/M,\mathbb{Z}/p\mathbb{Z})=\underline{H}^{n}_{\mathfrak{A}}(G/M,\mathbb{Z}/p\mathbb{Z})$. Applying Quillen’s theorem to $e$ we get that there is a positive integer $m$ and an $f\in H^{n}(G/M,\mathbb{Z}/p\mathbb{Z})$ such that $e^{p^{m}}=q_{G/M}(f)$. Let $\phi_{M}\colon\underline{H}^{n}_{\mathfrak{A}}(G/M,\mathbb{Z}/p\mathbb{Z})\to\underline{H}^{n}_{\mathfrak{A}}(G,M,\mathbb{Z}/p\mathbb{Z})$ be the map induced by the inclusion functor $\mathfrak{A}(G,M)\to\mathfrak{A}(G/M)$ as above. Then $\mathfrak{a}_{M}(\phi_{M}(e))=\mathfrak{a}_{N}(\underline{d})$, and hence
@@
$\pi_{i}\colon H^{i}(G,\mathbb{Z}/2\mathbb{Z})\to C(\mathcal{X}^{*}(G),\mathbb{Z}/2\mathbb{Z})$
@@
Let $C^{\bullet}(\mathcal{X}^{*}(G),\mathbb{Z}/2\mathbb{Z})$ denote the graded Boolean algebra associated to the Boolean ring $C(\mathcal{X}^{*}(G),\mathbb{Z}/2\mathbb{Z})$. By Proposition 7.11, the conditions of Proposition 7.17 apply to $G$. Hence, by Theorem 7.21, there is an $F$-isomorphism:
$\pi_{*}\colon H^{\bullet}(G,\mathbb{Z}/2\mathbb{Z})\to C^{\bullet}(\mathcal{X}^{*}(G),\mathbb{Z}/2\mathbb{Z}).$
This map is zero on the nilradical of $H^{\bullet}(G,\mathbb{Z}/2\mathbb{Z})$, so it induces an $F$-isomorphism:
(7.24.1) $H^{\bullet}(G,\mathbb{Z}/2\mathbb{Z})/\mathcal{N}(H^{\bullet}(G,\mathbb{Z}/2\mathbb{Z}))\to C^{\bullet}(\mathcal{X}^{*}(G),\mathbb{Z}/2\mathbb{Z})$
of graded Boolean algebras by Lemma 7.7, which must be an isomorphism by Lemma 7.13. Since the nilradical of $H^{\bullet}(G,\mathbb{Z}/2\mathbb{Z})$ consists of degree one elements, the claim follows. ∎
@@ 8. Proof of the main theorem and some consequences
The implication $(i)\Rightarrow(ii)$ is Theorem 6.6, while the implication $(ii)\Rightarrow(i)$ is Theorem 6.18. We already saw that $(ii)$ trivially implies $(iii)$ in the proof of Theorem 5.4. Now assume that $G$ is the maximal pro-$2$ quotient of a real projective profinite group $H$. As explained in [15, Section 10], $H^{\bullet}(H,\mathbb{Z}/2\mathbb{Z})$ is the connected sum of a dual algebra and a graded Boolean algebra. The pull-back map $H^{\bullet}(H,\mathbb{Z}/2\mathbb{Z})\to H^{\bullet}(G,\mathbb{Z}/2\mathbb{Z})$ is an isomorphism by the Rost–Voevodsky norm residue theorem [23], so we get that $G$ is cohomologically quasi-Boolean. Therefore the implication $(iii)\Rightarrow(iv)$ holds. To finish the proof of Theorem 1.7, we show the remaining implication $(iv)\Rightarrow(i)$ in the following
@@
By part $(i)$ of Proposition 7.11, $G$ has an open subgroup without $2$-torsion. By Remark 1.2, it will therefore be sufficient to show that every real embedding problem for $G$ has a solution. By Proposition 5.2, we need to show that any real $2$-embedding problem for $G$ has a solution. In fact we will prove something stronger.
@@
For every group $G$, let $G^{\#}$ denote the set of its conjugacy classes. For every $x\in G$, let $x^{\#}\in G^{\#}$ denote the conjugacy class of $x$, and, for every homomorphism $h\colon G\to H$ of groups, let $h^{\#}\colon G^{\#}\to H^{\#}$ denote the map on conjugacy classes induced by $h$. Now let $G$ be a profinite group. An embedding problem with lifting data $(\mathbf{E},f)$ for $G$ is an embedding problem $\mathbf{E}$:
@@
$\chi^{\bullet}(o(\mathbf{E}))=o(\mathbf{E}(\chi)),$
where $\chi^{\bullet}\colon H^{\bullet}(G,\mathrm{Ker}(\mathbf{E}))\to H^{\bullet}(H,\mathrm{Ker}(\mathbf{E}))$ is the pull-back map on cohomology.
@@
Let $G$ be a quasi-Boolean pro-$2$ group $G$. Then every $2$-embedding problem with lifting data for $G$ and with a kernel isomorphic to $\mathbb{Z}/2\mathbb{Z}$ has a solution.
@@
Let $(\mathbf{E},f)$ be an embedding problem with lifting data for $G$ as in Definition 8.2, and assume that its kernel is isomorphic to $\mathbb{Z}/2\mathbb{Z}$. Since the automorphism group of the latter is trivial, we get that $\mathbf{E}$ is central. Because $\mathbf{E}$ is equipped with lifting data, it is real, that is, for every subgroup $H\subseteq G$ of order $2$ the embedding problem $\mathbf{E}(i_{H})$ has a solution, where $i_{H}\colon H\to G$ is the inclusion map. Therefore, by Remark 8.5, the image of $o(\mathbf{E})$ under the homomorphism:
$\pi_{2}\colon H^{2}(G,\mathbb{Z}/2\mathbb{Z})\to C(\mathcal{X}^{*}(G),\mathbb{Z}/2\mathbb{Z})$
@@
Let $s\colon G\to B$ be such a solution. Let $r\colon\mathcal{X}^{*}(G)\to\mathbb{Z}/2\mathbb{Z}=\mathrm{Ker}(\alpha)$ be the map given by the rule:
@@
Since $s^{\#}\times f\colon\mathcal{X}^{*}(G)\to B^{\#}\times B^{\#}$ is continuous with finite image, and $r(x)$ only depends on $s^{\#}(x)$ and $f(x)$ for every $x\in\mathcal{X}^{*}(G)$, we get that $r$ is continuous. Therefore, by Corollary 7.24, there is a continuous homomorphism $\chi\colon G\to\mathbb{Z}/2\mathbb{Z}=\mathrm{Ker}(\alpha)$ whose image under the homomorphism
$\pi_{1}\colon H^{1}(G,\mathbb{Z}/2\mathbb{Z})\to C(\mathcal{X}^{*}(G),\mathbb{Z}/2\mathbb{Z})$
@@
using that $s$ and $\chi$ are homomorphisms and $\mathrm{Ker}(\alpha)$ is central. Therefore $\widetilde{\phi}$ is a homomorphism. Since $\alpha\circ\widetilde{\phi}=\alpha\circ s=\phi$, we get that $\widetilde{\phi}$ is a solution to $\mathbf{E}$. Now let $y\in\mathcal{Y}^{*}(G)$ be arbitrary. If $s^{\#}(y^{\#})=f(y^{\#})$, then $\widetilde{\phi}(y)=s(y)$, and hence $\widetilde{\phi}(y)^{\#}=s(y)^{\#}=s^{\#}(y^{\#})=f(y^{\#})$. If $s^{\#}(y^{\#})\neq f(y^{\#})$, then $\widetilde{\phi}(y)\neq s(y)$, so $\widetilde{\phi}(y)$ is the unique element of $\alpha^{-1}(\phi(y))$ distinct from $s(y)$. Since $\alpha^{-1}(\phi(y))$ contains an element of $f(y^{\#})$, as $\alpha$ is surjective, and this element is not $s(y)$, it must be $\widetilde{\phi}(y)$. So $\widetilde{\phi}(y)^{\#}=f(y^{\#})$ in this case, too. ∎
@@
such that the kernel of the quotient map $\pi_{k}\colon B/N_{k}\to B/N_{k+1}$ is isomorphic to $\mathbb{Z}/2\mathbb{Z}$ for every $k=0,1,\ldots,n-1$. Let $q_{k}\colon B\to B/N_{k}$ be the quotient map. Note that it will be sufficient to show that for every continuous homomorphism $h\colon G\to B/N_{k+1}$ such that $h^{\#}|_{\mathcal{X}^{*}(G)}=q_{k+1}^{\#}\circ f=\pi_{k}^{\#}\circ q_{k}^{\#}\circ f$ the embedding problem $\mathbf{E}_{k}$:
@@
with lifting data $q_{k}^{\#}\circ f$ has a solution for every $k=0,1,\ldots,n$. Indeed let $r_{k}\colon B/N_{k}\to A=B/N_{n}$ be the quotient map. Then we would get by descending induction on the index $k$ that the embedding problem:
@@
with lifting data $q_{k}^{\#}\circ f$ has a solution. The claim is now clear from the case $k=0$. However, $\mathbf{E}_{k}$ has a kernel isomorphic to $\mathbb{Z}/2\mathbb{Z}$, so $(\mathbf{E}_{k},q_{k}^{\#}\circ f)$ has a solution by Lemma 8.6. ∎
@@
First assume that $G$ is Boolean, i.e., $G$ is isomorphic to $\mathbb{B}(X)$ for some profinite space $X$. Then it is cohomologically quasi-Boolean by Theorem 1.7, so $H^{\bullet}(G,\mathbb{Z}/2\mathbb{Z})$ is the connected sum of a dual algebra $D^{\bullet}$ and a graded Boolean algebra $B^{\bullet}$. Assume that $D^{\bullet}$ is non-trivial, so there is a non-zero $c\in D^{1}\subset\textrm{Hom}(G,\mathbb{Z}/2\mathbb{Z})$. Then $c\cup c=0$ as $D^{2}=0$, so the restriction of $c$ onto every involution in the image of $i_{X}\colon X\to\mathcal{Y}(\mathbb{B}(X))$ is zero. Therefore, by the universal property of $\mathbb{B}(X)$ of Remark 6.3, the homomorphism corresponding to $c$ is also zero, a contradiction. Hence $G$ is cohomologically Boolean.
Next assume that $G$ is cohomologically Boolean. Then it is quasi-Boolean by Theorem 1.7, so $G$ is the free product of a free pro-$2$ group $F$ and a Boolean group $\mathbb{B}(X)$ for a profinite space $X$. Assume that $F$ is non-trivial. Then there is a non-zero $c\in\textrm{Hom}(F,\mathbb{Z}/2\mathbb{Z})$. Let $\pi_{1}\colon G\to F$ be the surjective homomorphism supplied by the universal property of free pro-$2$ products, and let $\overline{c}$ be the composition of $\pi_{1}$ and $c$. Then $c\cup c=0$, and hence $\overline{c}\cup\overline{c}=0$. Since $H^{\bullet}(G,\mathbb{Z}/2\mathbb{Z})$ is a graded Boolean algebra, this implies that $\overline{c}$ is zero, a contradiction. Therefore $G$ is Boolean. ∎
@@
Let $G$ be a Boolean group and let $f\colon G\to\mathbb{Z}/2\mathbb{Z}$ be a quasi-canonical homomorphism. By Remark 7.9, the homomorphism $f$ is characterised by the property that, for every $n>0$ and $c\in H^{n}(G,\mathbb{Z}/2\mathbb{Z})$, we have $c^{2}=c\cup k^{n}$. But $H^{\bullet}(G,\mathbb{Z}/2\mathbb{Z})$ is a graded Boolean algebra by Corollary 1.8, and hence $f$ is unique. Therefore, it is justified to call it the canonical homomorphism of the Boolean group $G$.
@@
By assumption, $G$ is isomorphic to $F(Y^{\prime})*_{2}\mathbb{B}(X^{\prime})$ for a set $Y^{\prime}$ and a profinite space $X^{\prime}$. We are actually going to show a stronger claim than in the theorem, namely, that the profinite spaces $X$ and $X^{\prime}$ are homeomorphic, and the sets $Y^{\prime}$ and $Y$ are bijective. By Theorem 6.11, the profinite spaces $X^{\prime}$ and $\mathcal{X}^{*}(G)$ are homeomorphic, and, as we saw in the proof of Corollary 7.24, the graded Boolean algebras $B^{\bullet}$ and $C^{\bullet}(\mathcal{X}^{*}(G),\mathbb{Z}/2\mathbb{Z})$ are isomorphic (see (7.24.1)). Therefore, by Stone duality of Theorem 2.11, the profinite spaces $X$ and $X^{\prime}$ are homeomorphic, too. Using Notation 6.17, we have $G_{*}=F(Y^{\prime})_{*}\oplus\mathbb{B}(X^{\prime})_{*}$. Hence we have $H^{1}(G,\mathbb{Z}/2\mathbb{Z})=H^{1}(F(Y^{\prime}),\mathbb{Z}/2\mathbb{Z})\oplus H^{1}(\mathbb{B}(X^{\prime}),\mathbb{Z}/2\mathbb{Z})$. However, since $H^{1}(\mathbb{B}(X^{\prime}),\mathbb{Z}/2\mathbb{Z})\cong B^{1}$ by the above, we get that $H^{1}(F(Y^{\prime}),\mathbb{Z}/2\mathbb{Z})\cong D^{1}$, and hence $H^{1}(G,\mathbb{Z}/2\mathbb{Z})=D^{1}\oplus B^{1}$. Therefore, $\mathbb{Z}/2\mathbb{Z}^{\oplus Y^{\prime}}\cong\mathbb{Z}/2\mathbb{Z}^{\oplus Y}$, and hence there is a bijection between the sets $Y^{\prime}$ and $Y$. ∎